# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ 3.2.2 Sub-optimal Calibration and Agreement With NR Waveforms
State-of-the-art waveform models are built by combining and resumming information from different analytical methods, such as PN approximation and gravitational self-force theory, and then calibrating/validating against NR simulations and merger-ringdown waveforms in the test-particle limit, which are obtained by solving the Teukolsky equation. The assessment of the accuracy of the waveform models from the two main waveform families (notably EOB and IMRPhenom models) can be found in Pompili et al. (2023); Ramos-Buades et al. (2023a); Nagar et al. (2023); Thompson et al. (2023); Mac Uilliam et al. (2024); Hu & Veitch (2022); Dhani et al. (2024). Due to the number of calibration parameters and the large number of NR simulations at disposal, it is especially important to devise a computationally efficient and flexible calibration procedure. For instance, in calibrating the SEOBNRv5HM model (Pompili et al., 2023), the authors quantified the agreement with NR waveforms in a Bayesian fashion and employed nested sampling to obtain posterior distributions for the calibration parameters. State-of-the-art waveform models use best-fit estimates across the physical parameter space for their calibration parameters. Providing instead a probability distribution, modeled for example through a multidimensional Gaussian mixture, would allow accounting for uncertainty estimates due to sub-optimal fits, and could mitigate waveform systematics at high SNR. Other proposed methods to marginalise over waveform modeling uncertainties include Gaussian process regression (Moore & Gair, 2014; Doctor et al., 2017; Williams et al., 2020; Khan, 2024), or introducing frequency-dependent amplitude and phase corrections, as in the case of detector calibration uncertainty (Read, 2023).
@@
Calibration parameters typically enter in waveform models as higher-order PN coefficients, which are currently unknown. Including higher-order analytical information, while pushing the calibration parameters at even higher orders, could improve the accuracy of current waveform models, but requires careful studies on how to incorporate and resum this information (Nagar & Rettegno, 2021; Pompili et al., 2023; Nagar et al., 2023) Nonetheless, neglecting higher-order PN terms carries an error which might become relevant with updates to current detectors and XG detectors, but could be mitigated by marginalizing over higher-order PN coefficients as new model parameters (Owen et al., 2023). Incorporating results from the post-Minkowskian (PM) approximation (Kosower et al., 2022; Bjerrum-Bohr et al., 2022; Antonelli et al., 2019; Di Vecchia et al., 2023), a weak fields expansion in $G$ at all orders in the velocity, is also promising, particularly for highly eccentric binaries for which relativistic velocities can be reached at each periastron passage even in the weak field regime. While PM results have not yet been incorporated in state-of-the-art waveform models for bound orbits, remarkable agreement has been obtained comparing PM-improved EOB models to NR for scattering orbits (Khalil et al., 2022; Damour & Rettegno, 2023; Rettegno et al., 2023; Buonanno et al., 2024).
@@
The calibration procedure imposes that the waveform model agrees, as much as possible and for the entire coalescence, with the NR waveform. This is often quantified by computing the unfaithfulness (or mismatch) $\mathcal{M}$ between the model and NR waveform. As detectors become more sensitive and the SNR increases, the accuracy requirements become more stringent, thus demanding smaller unfaithfulness values. Accuracy requirements are usually formulated in terms of an indistinguishability criterion (Flanagan & Hughes, 1998; Lindblom et al., 2008; McWilliams et al., 2010; Chatziioannou et al., 2017; Pürrer & Haster, 2020), which states that if two waveforms fulfill the condition
@@
for a given power spectral density (PSD) and SNR, then these waveforms are considered indistinguishable, and differences in the recovered parameters are expected to be smaller than statistical errors. Here $D$ is an unknown coefficient, usually set to the number of intrinsic parameters of the source (Chatziioannou et al., 2017) or tuned with synthetic injections at increasing SNR (Pürrer & Haster, 2020). Being sufficient, but not necessary, this criterion is generally too conservative, and, if it is violated, differences are not necessarily measurable, or may appear in a subset of parameters in which one is not typically interested (Pürrer & Haster, 2020; Ossokine et al., 2020). Toubiana & Gair (2024) recently proposed a correction to the standard indistinguishability criterion by revisiting some of the hypotheses under which it is derived, and employed it to quantify apparent deviations from GR due to waveform inaccuracies (Toubiana et al., 2023).
The state-of-the-art multipolar, aligned-spin SEOBNRv5HM model, which has median unfaithfulness of $1.01\times 10^{-3}$ against 442 NR waveforms (when using the O5 PSD (Barsotti et al., 2018), maximizing over the total binary mass in the range $[20-300]M_{\odot}$), would lead to a false deviation from GR in measuring the QNM (complex) frequencies of a heavy massive mass ratio 2 binary black hole when observed in LISA with an SNR $\mathcal{O}(100)$ (Toubiana et al., 2023). This issue occurs because for such massive binary black holes, the majority of the SNR lies in the merger-ringdown stage. By contrast, a stellar-mass binary black hole with mass ratio 6, observable in O5, would not incorrectly lead to a violation of GR at SNR $75$ (Ghosh et al., 2021), because in this case a large portion of the SNR is accumulated during the inspiral stage. Normally, the accuracy of waveform models gets worse toward merger, where the presence of higher-order modes becomes more and more important, while their modeling is quite challenging. The recent study of Kapil et al. (2024) investigated the impact of inference biases from sub-optimal waveform calibration on a realistic population of binary black holes in XG detectors. They considered two quasi-circular, nonprecessing waveform models of the same family (namely, IMRPhenomD (Khan et al., 2016) and IMRPhenomXAS (Pratten et al., 2020)) and estimated a mismatch requirement of $\sim 10^{-5}$ for $99\%$ of the events with $\rm{SNR}>100$ not to be biased.
Inaccuracies in NR waveforms, due to, e.g., numerical truncation errors and issues with GW extraction and extrapolation, are typically at least one order of magnitude smaller than errors between semi-analytic models and NR (Pürrer & Haster, 2020). Nonetheless, they are expected to become relevant with updates to current detectors and XG detectors, especially for binaries with asymmetric masses and orbits inclined with respect to the line of sight (Pürrer & Haster, 2020; Ferguson et al., 2021; Jan et al., 2023).
@@ 4.1 Gravitational Lensing
As GW detectors get upgraded and new ones join the network, more and more distant mergers can be observed. This increases the chance of having a matter density crossing the GW travel path, possibly leading to gravitational lensing. Depending on the lens properties and the lens-source geometry, different effects can be observed. For the best-aligned and most massive cases, we are in the geometric optics limit and lensing leads to several copies or “images” of the initial signal. These images have the same frequency evolution but are delayed in time, (de)magnified, and can undergo an overall phase shift. When the time delay is large enough, these images are distinct, and we face strong lensing (Takahashi & Nakamura, 2003; Dai & Venumadhav, 2017). For ground-based detectors, typical lenses are galaxies and galaxy clusters (Schneider et al., 1992). For smaller time delays, corresponding to less aligned systems and lighter lenses, one has millilensing, where the various images overlap and sum to a non-trivial signal in-band (Liu et al., 2023). This is expected to be due to heavy black holes, or dark matter over-densities, for example. Finally, when the GW wavelength is comparable to or greater than the size of the lens, we need to perform the full wave-optics treatment (Takahashi & Nakamura, 2003), and lensing leads to frequency-dependent beating patterns known as microlensing. For ground-based detectors, typical lens sources are individual stars, black holes, or dark-matter overdensities (Wright & Hendry, 2021). It is also important to note there can be interplay between these different types of lensing. When strong lensing happens, one or more of the images may undergo micro or millilensing because of individual objects present in the strong lens (Seo et al., 2022; Mishra et al., 2021; Meena et al., 2022).
@@
False GR deviations could be expected when GR signals are distorted. For strong lensing, one can have such an effect for specific values of the overall phase shift. In particular, it can take only three distinct values: 00, $\pi/2$, or $\pi$, corresponding to a minimum, saddle point, or maximum of the Fermat potential, and referred to as Type I, II, and III images, respectively (Dai & Venumadhav, 2017; Ezquiaga et al., 2021). Under all circumstances, Type I and III images are indistinguishable for the GR case because they correspond to no shift or a sign flip in the polarization, which cannot be detected (Ezquiaga et al., 2021). For Type II images, on the other hand, detectability is possible when the GW displays higher-order modes. In this case, the phase has different pre-factors for different frequency modes and is not degenerate with the (frequency independent) lensing phase shift anymore (Ezquiaga et al., 2021). This can be used to detect strong lensing based on a single image, although it requires rather large SNRs and very asymmetric, precessing or eccentric systems (Wang et al., 2021; Janquart et al., 2021; Ezquiaga et al., 2021; Vijaykumar et al., 2023a). When analyzing Type II images under the unlensed assumptions, one can face losses in SNR, possibly missing the event with template searches (Wang et al., 2021), or biases in parameter estimation (Janquart et al., 2021; Vijaykumar et al., 2023a). Therefore, one can expect this non-trivial feature to also be picked up when searching for GR deviations. For example, this is the case with modified dispersion relations that change the frequency evolution of the GW phase in a way possibly similar to lensing (Ezquiaga et al., 2022). The link between Type II images and GR deviations is also highlighted in Wright et al. (2024), where the authors show that some GR deviations are flagged by Type II search pipelines.
@@
The cases of millilensing and microlensing are even more favorable in leading to spurious GR deviations being detected since they both lead to a non-trivial signal in the detection band, although the nature of the resulting image is different between the two cases (Takahashi & Nakamura, 2003; Wright & Hendry, 2021; Liu et al., 2023). When analyzing such signals with traditional GR templates, one expects imperfect modeling of the signal, leading to coherent power left in the data (Janquart et al., 2023). This is also confirmed in Mishra et al. (2023) for some tests of GR. In this study, the authors show that milli and microlensed signals can lead to spurious deviations from GR, sometimes with a high significance. However, it is also important to note that adapted lensing pipelines also clearly see these events as being lensed. Therefore, the GR deviation would probably not be confirmed as it would be explained via lensing, underlying the importance of accounting for possible astrophysical effects on the GW signals when looking for GR deviations. The link between GR deviations and micro and millilensing is also further confirmed in Wright et al. (2024), where the authors show that some deviations of GR lead to false positives in micro and millilensing searches. In the case of a multi-messenger lensing event in which the GW lensed signal is in the wave optics regime but the electromagnetic signal is in geometric optics (which is to be expected given their higher frequency), the speed of propagation of GWs could appear to be superluminal due to the waveform distortions (Ezquiaga et al., 2020), although no information actually arrives faster than light (Suyama, 2020).
@@
A crucial approximation in these studies is the exclusion of the effect of parallel-transporting the polarization tensor across the lensing geometry and the treatment of GWs as scalar waves which become increasingly violated as one moves from the weak gravity limit. Recent studies (Harte, 2019; Cusin & Lagos, 2020) have pointed out the consequences of such an approximation and started treating GWs as a tensor field. It is pointed out that there is no notion of a unique “propagation direction” as can be defined in the geometric optics limit as well as the wave optics treatment for a scalar wave. Similarly, strong gravity effects could add extra phenomenology (Yin & He, 2024).
@@ 4.2 Environmental Effects
The current waveform models can be referred to as vacuum templates as they only describe GWs from isolated binary systems in a vacuum environment, neglecting realistic astrophysical surroundings of the source. However, in reality, the binary is always in an astrophysical environment that impacts the binary’s orbital evolution and hence results in a GW signal from the binary different than the vacuum template. There are many scenarios in which the GW signal from an environment-embedded binary system could be different from its corresponding vacuum signal. These are, but not limited to, (i) the source resides in a dense environment (Ostriker, 1999; Kim & Kim, 2007; Kim et al., 2008; Barausse, 2007) such as dense cores of massive stars (Loeb, 2016; Fedrow et al., 2017; D’Orazio & Loeb, 2018), accretion disks of active galactic nuclei (McKernan et al., 2012; Barausse et al., 2014; Bartos et al., 2017; Stone et al., 2017; Caputo et al., 2020; Toubiana et al., 2021; Sberna et al., 2022b), and star clusters (see, e.g., Mandel & Broekgaarden (2022)), (ii) the source resides in a dark matter halo (Eda et al., 2013; Coogan et al., 2022; Gondolo & Silk, 1999; Bertone & Merritt, 2005; Barausse et al., 2014; Eroshenko, 2016; Boucenna et al., 2018), and (iii) the source is immersed in a strong electromagnetic field (Punsly, 1998a, b). Moreover, the peculiar acceleration of the source with respect to the observer, i.e., time-varying Doppler shift (Bonvin et al., 2017; Tamanini et al., 2020; Inayoshi et al., 2017; Chen et al., 2019) and the acceleration of the universe, i.e., time-varying redshift itself (Seto et al., 2001; Nishizawa et al., 2012; Bonvin et al., 2017) could lead to GW signals being different from vacuum templates.
@@
The detailed modeling of different environmental effects on the binary’s GW signal is challenging and requires computationally expensive NR simulations (Fedrow et al., 2017). However, in the literature, these effects have been approximated as a correction to the vacuum GW signal’s PN phase evolution. For example, at the leading order, dynamical friction due to gas accretion can be modeled as a $-5.5$PN correction whereas collisionless (collisional) accretion can be modeled as a $-4.5$PN ($-5.5$PN) correction (Eddington, 1988; Begelman, 1977; Barausse, 2007; Cardoso & Maselli, 2020). The accretion and dynamical friction due to a scalar dark matter cloud give rise to a $-4$PN and $-5.5$PN correction, respectively, to the phase at the leading order (Boudon et al., 2023). Electromagnetic effects have been computed at next-to-leading order (at 3PN) by taking into account the whole electromagnetic structure of a star. The leading magnetic corrections at 2PN order (assuming a constant and aligned magnetic dipole) to the GW phase are found to be comparable to a 1.5PN point-particle effect (Henry et al., 2023a, b). Phase correction due to the line-of-sight peculiar acceleration of the source has been computed up to $3.5$PN order (Tamanini et al., 2020; Vijaykumar et al., 2023b) while the acceleration of the universe leads to a $-4$PN correction to the phase at leading order (Seto et al., 2001; Nishizawa et al., 2012).
@@
It has been argued that the magnitude of the environmental (Barausse et al., 2014; Chen & Shen, 2019) and cosmological (Bonvin et al., 2017) effects are expected to be quite small and hence could be neglected for ground-based detectors. However, there could be scenarios where these effects are non-negligible, e.g., stellar-mass compact binaries would merge around a supermassive black hole and one can still get a significant deviation from the vacuum template in the bands of LIGO/Virgo/KAGRA detectors (Vijaykumar et al., 2023b). Moreover, near supermassive black holes, in galactic nuclei, triple systems of stars are common and they mostly are hierarchical in nature (Samsing & Ilan, 2018, 2019; Tory et al., 2022), i.e., a tight inner binary is orbiting a tertiary on a wider orbit which forms the outer binary. In these hierarchical triples, the tertiary brings interesting features to the GW signal emitted by the inner binary, e.g., the oscillation of eccentricity and inclination of the inner binary’s orbit due to the Kozai-Lidov mechanism (Kozai, 1962; Lidov, 1962). Such oscillations could modify the frequency evolution of the inner binary and this needs to be taken into account in waveform modeling (Gupta et al., 2020; Chandramouli & Yunes, 2022).
@@
A recent study by Santoro et al. (2023) showed that particularly large environmental effects can significantly bias the parameter estimation if vacuum templates are used for the analysis, even when not directly detectable by LIGO-like instruments. Although this bias requires extremely dense environments that are not predicted by standard astrophysical models, it would be important to find out if such biases in parameters could lead to false GR violations for more sensitive XG detectors.
@@
Likewise, ringdown templates are simple and based on predictions from vacuum GR. Modifications of GR usually lead to extra polarizations or include degrees of freedom with different modes, introducing a simple handle to test for beyond-GR physics. However, environmental effects, such as accretion disks, dark matter halos or any form of matter outside of black holes introduces low-frequency modes or drastic changes to higher overtones, de-stabilizing the spectrum (Barausse et al., 2014; Jaramillo et al., 2022; Cheung et al., 2022). Concrete examples suggest that spectral instability of the dominant mode introduces changes in the waveform only well after coalescence, but the relevance of overtone instability for time-domain waveforms still needs to be well understood (Cardoso et al., 2024).

### v2

@@ 3.2.2 Sub-optimal Calibration and Agreement With NR Waveforms
State-of-the-art waveform models are built by combining and resumming information from different analytical methods, such as PN approximation and gravitational self-force theory, and then calibrating/validating against NR simulations and merger-ringdown waveforms in the test-particle limit, which are obtained by solving the Teukolsky equation. The assessment of the accuracy of the waveform models from the two main waveform families (notably EOB and IMRPhenom models) can be found in [320, 322, 335, 315, 336, 337, 248]. Due to the number of calibration parameters and the large number of NR simulations at disposal, it is especially important to devise a computationally efficient and flexible calibration procedure. For instance, in calibrating the SEOBNRv5HM model [320], the authors quantified the agreement with NR waveforms in a Bayesian fashion and employed nested sampling to obtain posterior distributions for the calibration parameters. State-of-the-art waveform models use best-fit estimates across the physical parameter space for their calibration parameters. Providing instead a probability distribution, modeled for example through a multidimensional Gaussian mixture, would allow accounting for uncertainty estimates due to sub-optimal fits, and could mitigate waveform systematics at high SNR [338]. Other proposed methods to marginalise over waveform modeling uncertainties include Gaussian process regression [339, 340, 341, 342], or introducing frequency-dependent amplitude and phase corrections, as in the case of detector calibration uncertainty [164]. While these methods may obscure small deviations from GR, particularly around the merger phase, significant deviations that exceed the estimated modeling uncertainties should still be detectable.
@@
Calibration parameters typically enter in waveform models as higher-order PN coefficients, which are currently unknown. Including higher-order analytical information, while pushing the calibration parameters at even higher orders, could improve the accuracy of current waveform models, but requires careful studies on how to incorporate and resum this information [343, 320, 335]. Nonetheless, neglecting higher-order PN terms carries an error which might become relevant with updates to current detectors and XG detectors, but could be mitigated by marginalizing over higher-order PN coefficients as new model parameters [344]. Incorporating results from the post-Minkowskian (PM) approximation [345, 346, 347, 348], a weak fields expansion in $G$ at all orders in the velocity, is also promising, particularly for highly eccentric binaries for which relativistic velocities can be reached at each periastron passage even in the weak field regime. While PM results have not yet been incorporated in state-of-the-art waveform models for bound orbits, remarkable agreement has been obtained comparing PM-improved EOB models to NR for scattering orbits [349, 350, 351, 352].
@@
The calibration procedure imposes that the waveform model agrees, as much as possible and for the entire coalescence, with the NR waveform. This is often quantified by computing the unfaithfulness (or mismatch) $\mathcal{M}$ between the model and NR waveform. As detectors become more sensitive and the SNR increases, the accuracy requirements become more stringent, thus demanding smaller unfaithfulness values. Accuracy requirements are usually formulated in terms of an indistinguishability criterion [353, 354, 355, 356, 357], which states that if two waveforms fulfill the condition
@@
for a given power spectral density (PSD) and SNR, then these waveforms are considered indistinguishable, and differences in the recovered parameters are expected to be smaller than statistical errors. Here $D$ is an unknown coefficient, usually set to the number of intrinsic parameters of the source [356] or tuned with synthetic GW signals at increasing SNR [357]. Being sufficient, but not necessary, this criterion is generally too conservative, and, if it is violated, differences are not necessarily measurable, or may appear in a subset of parameters in which one is not typically interested [357, 358]. Toubiana & Gair [359] recently proposed a correction to the standard indistinguishability criterion by revisiting some of the hypotheses under which it is derived, and employed it to quantify apparent deviations from GR due to waveform inaccuracies [360].
The state-of-the-art multipolar, aligned-spin SEOBNRv5HM model shows a median unfaithfulness of $1.01\times 10^{-3}$ against 442 NR waveforms when using the O5 PSD [361] and considering binary total masses in the range $[20-300]M_{\odot}$. Using this model would lead to a false deviation from GR when measuring the QNM (complex) frequencies of a massive binary black hole with a mass ratio of 2, as observed in LISA with an SNR $\mathcal{O}(100)$ [360]. This issue occurs because for such massive binary black holes, the majority of the SNR lies in the merger-ringdown stage. By contrast, a stellar-mass binary black hole with mass ratio 6, observable in O5, would not incorrectly lead to a violation of GR at SNR $75$ [282], because in this case a large portion of the SNR is accumulated during the inspiral stage. Normally, the accuracy of waveform models gets worse toward merger, where the presence of higher-order modes becomes more and more important, while their modeling is quite challenging. The recent study of [362] investigated the impact of inference biases from sub-optimal waveform calibration on a realistic population of binary black holes in XG detectors. They considered two quasi-circular, nonprecessing waveform models of the same family (namely, IMRPhenomD [363] and IMRPhenomXAS [364]) and estimated a mismatch requirement of $\sim 10^{-5}$ for $99\%$ of the events with $\rm{SNR}>100$ not to be biased.
Inaccuracies in NR waveforms, due to, e.g., numerical truncation errors and issues with GW extraction and extrapolation, are typically at least one order of magnitude smaller than errors between semi-analytic models and NR [357]. Nonetheless, they are expected to become relevant with updates to current detectors and XG detectors, especially for binaries with asymmetric masses and orbits inclined with respect to the line of sight [357, 365, 366].
@@ 4.1 Gravitational Lensing
As GW detectors get upgraded and new ones join the network, more and more distant mergers can be observed. This increases the chance of having a matter density crossing the GW travel path, possibly leading to gravitational lensing. Depending on the lens properties and the lens-source geometry, different effects can be observed. For the best-aligned and most massive cases, we are in the geometric optics limit and lensing leads to several copies or “images” of the initial signal. These images have the same frequency evolution but are delayed in time, (de)magnified, and can undergo an overall phase shift. When the time delay is large enough, these images are distinct, and we face strong lensing [367, 368]. For ground-based detectors, typical lenses are galaxies and galaxy clusters [369]. For smaller time delays, corresponding to less aligned systems and lighter lenses, one has millilensing, where the various images overlap and sum to a non-trivial signal in-band [370]. This is expected to be due to heavy black holes, or dark matter over-densities, for example. Finally, when the GW wavelength is comparable to or greater than the size of the lens, we need to perform the full wave-optics treatment [367], and lensing leads to frequency-dependent beating patterns known as microlensing. For ground-based detectors, typical lens sources are individual stars, black holes, or dark-matter overdensities [371]. It is also important to note there can be interplay between these different types of lensing. When strong lensing happens, one or more of the images may undergo micro or millilensing because of individual objects present in the strong lens [372, 373, 374].
@@
False GR deviations could be expected when GR signals are distorted. For strong lensing, one can have such an effect for specific values of the overall phase shift. In particular, it can take only three distinct values: 00, $\pi/2$, or $\pi$, corresponding to a minimum, saddle point, or maximum of the Fermat potential, and referred to as Type I, II, and III images, respectively [368, 375]. Under all circumstances, Type I and III images are indistinguishable from the GR case because they correspond to no shift or a sign flip in the polarization, which cannot be detected [375]. For Type II images, on the other hand, detectability is possible when the GW displays higher-order modes. In this case, the phase has different pre-factors for different frequency modes and is not degenerate with the (frequency independent) lensing phase shift anymore [375]. This can be used to detect strong lensing based on a single image, although it requires rather large SNRs and very asymmetric, precessing or eccentric systems [376, 377, 375, 378]. When analyzing Type II images under the unlensed assumptions, one can face losses in SNR, possibly missing the event with template searches [376], or biases in parameter estimation [377, 378]. Therefore, one can expect this non-trivial feature to also be picked up when searching for GR deviations [379]. For example, this is the case with modified dispersion relations that change the frequency evolution of the GW phase in a way possibly similar to lensing [380]. The link between Type II images and GR deviations is also highlighted in [381], where the authors show that some GR deviations are flagged by Type II lensing search pipelines.
@@
The cases of millilensing and microlensing are even more favorable in leading to spurious GR deviations being detected since they both lead to a non-trivial signal in the detection band, although the nature of the resulting image is different between the two cases [367, 371, 370]. When analyzing such signals with traditional GR templates, one expects imperfect modeling of the signal, leading to coherent power left in the data [382]. This is also confirmed in [383] for some tests of GR. In this study, the authors show that milli and microlensed signals can lead to spurious deviations from GR, sometimes with a high significance. However, it is also important to note that adapted lensing pipelines also clearly see these events as being lensed. Therefore, the GR deviation would probably not be confirmed as it would be explained via lensing, underlying the importance of accounting for possible astrophysical effects on the GW signals when looking for GR deviations. The link between GR deviations and micro and millilensing is also further confirmed in [381], where the authors show that some deviations of GR lead to false positives in micro and millilensing searches. In the case of a multi-messenger lensing event in which the GW lensed signal is in the wave optics regime but the electromagnetic signal is in geometric optics (which is to be expected given their higher frequency), the speed of propagation of GWs could appear to be superluminal due to the waveform distortions [384], although no information actually arrives faster than light [385].
@@
A crucial approximation in these studies is the exclusion of the effect of parallel-transporting the polarization tensor across the lensing geometry and the treatment of GWs as scalar waves which become increasingly violated as one moves from the weak gravity limit. Recent studies [386, 387] have pointed out the consequences of such an approximation and started treating GWs as a tensor field. It is pointed out that there is no notion of a unique “propagation direction” as can be defined in the geometric optics limit as well as the wave optics treatment for a scalar wave. Similarly, strong gravity effects could add extra phenomenology [388].
@@ 4.2 Environmental Effects
The current waveform models can be referred to as vacuum templates as they only describe GWs from isolated binary systems in a vacuum environment, neglecting realistic astrophysical surroundings of the source. However, in reality, the binary is always in an astrophysical environment that impacts the binary’s orbital evolution and hence results in a GW signal from the binary different than the vacuum template. There are many scenarios in which the GW signal from an environment-embedded binary system could be different from its corresponding vacuum signal. These are, but not limited to, (i) the source resides in a dense environment [389, 390, 391, 392] such as dense cores of massive stars [393, 394, 395], accretion disks of active galactic nuclei [396, 32, 397, 398, 399, 400, 401], and star clusters (see, e.g., [402]), (ii) the source resides in a dark matter halo [403, 404, 405, 406, 32, 407, 408], and (iii) the source is immersed in a strong electromagnetic field [409, 410]. Moreover, the peculiar acceleration of the source with respect to the observer, i.e., time-varying Doppler shift [411, 412, 413, 414] and the acceleration of the universe, i.e., time-varying redshift [415, 416, 411] itself could lead to GW signals being different from vacuum templates.
The situation where there is a massive bosonic field that is amplified around black holes via superradiance (see, e.g., [417]) is also sometimes considered an environmental effect and can similarly lead to deviations from a vacuum binary black hole signal. However, there are significant differences in this case compared to the environmental effects considered above. Most importantly, in this case the size of the deviation is set by the universal properties of the boson and the properties of the binary, not the specifics of where the binary formed. This makes the deviations more similar to a deviation from GR (which also depends on the properties of the binary and some universal parameters). However, there are many ways to distinguish binaries of black holes with boson clouds from GR deviations. Some of these are discussed in Sec. 4.3.2, since the emitted GW signal in such a scenario will be similar to the one from binaries of black hole mimickers (e.g., there will be tidal effects from the boson clouds). Additionally, since the superradiant growth of the clouds is only possible for certain pairs of black hole masses and spins (see, e.g., [418]), it should be easy to distinguish this case from modified gravity (or black hole mimickers) when considering the population. The time dependence of the tidal deformability and non-black hole multipole moments due to perturbations or even disruption of the clouds due to the effects of the other black hole (see, e.g., [419, 420, 421]) should allow one to distinguish the boson cloud case from black hole mimickers even for individual sources. Additionally, one can obtain constraints on the boson mass from the contributions from the superradiant instability to the stochastic background of GWs [422, 423]. Furthermore, boson clouds are expected to emit a nearly periodic and long-duration GW signal [422, 423] and no evidence of such signals is found in current GW data, which provides constraints on the ultra-light scalar boson field mass (see, e.g., [424, 425, 426, 427]).
@@
Returning to environmental effects proper, the detailed modeling of different environmental effects on the binary’s GW signal is challenging and requires computationally expensive NR simulations [394]. However, in the literature, these effects have been approximated as a correction to the vacuum GW signal’s PN phase evolution. For example, at the leading order, dynamical friction due to gas accretion can be modeled as a $-5.5$PN correction whereas collisionless (collisional) accretion can be modeled as a $-4.5$PN ($-5.5$PN) correction [428, 429, 392, 430]. The accretion and dynamical friction due to a scalar dark matter cloud give rise to a $-4$PN and $-5.5$PN correction, respectively, to the phase at the leading order [431]. Electromagnetic effects have been computed at next-to-leading order (at 3PN) by taking into account the whole electromagnetic structure of a star. The leading magnetic corrections at 2PN order (assuming a constant and aligned magnetic dipole) to the GW phase are found to be comparable to a 1.5PN point-particle effect [432, 433]. Phase correction due to the line-of-sight peculiar acceleration of the source has been computed up to $3.5$PN order [412, 434] while the acceleration of the universe leads to a $-4$PN correction to the phase at leading order [415, 416].
@@
It has been argued that the magnitude of the environmental [32, 435] and cosmological [411] effects are expected to be quite small and hence could be neglected for ground-based detectors. However, there could be scenarios where these effects are non-negligible, e.g., stellar-mass compact binaries would merge around a supermassive black hole and one can still get a significant deviation from the vacuum template in the bands of LIGO/Virgo/KAGRA detectors [434]. Moreover, near supermassive black holes, in galactic nuclei, triple systems of stars are common and they mostly are hierarchical in nature [436, 437, 438], i.e., a tight inner binary is orbiting a tertiary on a wider orbit which forms the outer binary. In these hierarchical triples, the tertiary brings interesting features to the GW signal emitted by the inner binary, e.g., the oscillation of eccentricity and inclination of the inner binary’s orbit due to the Kozai-Lidov mechanism [439, 440]. Such oscillations could modify the frequency evolution of the inner binary and this needs to be taken into account in waveform modeling [441, 442].
@@
A recent study by Santoro et al. [443] showed that particularly large environmental effects can significantly bias the parameter estimation if vacuum templates are used for the analysis, even when not directly detectable by LIGO-like instruments. Although this bias requires extremely dense environments that are not predicted by standard astrophysical models, it would be important to find out if such biases in parameters could lead to false GR violations for more sensitive XG detectors.
@@
Likewise, ringdown templates are simple and based on predictions from vacuum GR. Modifications of GR usually lead to extra polarizations or include degrees of freedom with different modes, introducing a simple handle to test for beyond-GR physics. However, environmental effects, such as accretion disks, dark matter halos or any form of matter outside of black holes introduces low-frequency modes or drastic changes to higher overtones, de-stabilizing the spectrum [32, 444, 445]. Concrete examples suggest that spectral instability of the dominant mode introduces changes in the waveform only well after coalescence, but the relevance of overtone instability for time-domain waveforms still needs to be well understood [446].